What will be the decrese in the value of g on going at a height R/2 above the earth's surface ?

Grade:11

1 Answers

Vikas TU
5 years ago
gravity acceleration above the earth is given by the formulae: => geff. = GM/(h + R)^2
where h is the distance from the earth’s surface.
geff. = GM/(R + R/2)^2 => 4GM/9R^2
 
at earth’s surface the g = GM/R^2
decrease in value of g => GM/R^2 –  4GM/9R^2 => 5GM/9R^2
Where G is gravity constant and M is the earth’s mass and R is the radius of earth.